# Reinforcement Learning Enhancing Entanglement for Two-Photon-Driven Rabi Model

### v3

@@ Abstract
A control scheme is proposed that leverages reinforcement learning to enhance entanglement by modulating the two-photon-driven amplitude in a Rabi model. The quantum phase diagram versus the amplitude of the two-photon process and the coupling between the cavity field and the atom in the Rabi model, is indicated by the energy spectrum of the hybrid system, the witness of entanglement, second order correlation, and negativity of Wigner function. From a dynamical perspective, the behavior of entanglement can reflect the phase transition and the reinforcement learning agent is employed to produce temporal sequences of control pulses to enhance the entanglement in the presence of dissipation. The entanglement can be enhanced in different parameter regimes and the control scheme exhibits robustness against dissipation. The replaceability of the controlled system and the reinforcement learning module demonstrates the generalization of this scheme. This research paves the way of positively enhancing quantum resources in non-equilibrium systems.
@@ I Introduction
The interaction between a two-level system and an electromagnetic field, fully described by the quantum Rabi model PR51652 ; RPP691325 is a platform which has widespread utility ranging from the study of fundamental physics Haroche2006 to the cutting edge of quantum information RMP91025005 . In the regime of cavity quantum electrodynamics (QED), when the cavity-qubit coupling rate is strong enough to dominate dissipation rates or reaches the domain of transition frequencies of the cavity or qubit, the counter-rotating term cannot be ignored, and the system is best described by the quantum Rabi model RMP91025005 ; PRL120-183601 . Once the coupling is strong enough, a wide range of interesting phenomena occurs, including innovative quantum phase transitions in the single-qubit-single-oscillator model when the frequency of the oscillator is much lower than that of the qubit PRA87-013826 ; PRL115180404 , generation of pure-cat-state PRA81042311 ; PRA96043834 , strongly entangled states, and novel steady-state squeezing PRL131223604 . To date, in specially designed architectures, the qubit-cavity coupling rate can reach the considerable fraction of the cavity transition frequency NP6772 . By inductively coupling a flux qubit and an LC oscillator via Josephson junctions, ultrastrong-coupling regime with unconventional transition spectra has been realized  NP1344 . Quantum simulation in circuit QED provides an experimental access to the physics of the ultrastrong-and deep-strong coupling regimes of an effective quantum Rabi model PRX2021007 ; NC8779 . The strong coupling limit of cavity QED has been demonstrated in a solid-state system by the vacuum Rabi mode splitting and the corresponding avoided crossings Nature431162 .
@@
Machine-learning techniques are emerging as effective tools in exploring physics Murphy2012 ; Sutton2018 ; Murphy2018 ; PRX8031084 , and among them, reinforcement learning (RL) has the potential to find optimal control fields to engineer system dynamics, given that the evolution of these systems is governed by specific differential equations. Time-varying control proposals has been employed to produce spin squeezing in a collective spin systems PRA103032601 ; AdP5362400056 . These works establishes a foundational framework to generate specific quantum states by reinforcement learning. Reinforcement learning can successfully steer the quantum Kapitza oscillator to a Floquet-engineered stable inverted position amidst noise, with applications in quantum information and quantum opticsPRB98224305 .
@@ II The model
where $\hat{a}^{\dagger}$($\hat{a})$ is the creation (annihilation) operator for the bosonic cavity mode and $\hat{\sigma}_{z}$ is the Pauli $z$-basis operator with commute relation $[\hat{\sigma}_{+},\hat{\sigma}_{-}]=\hat{\sigma}_{z}$. $\delta_{c(q)}=\omega_{c(q)}-\omega_{p}$, where $\omega_{c}$, $\omega_{q}$ and $\omega_{p}$ are the frequencies of the cavity, qubit and parametrical driving, respectively. The parameter $g$ is the coupling strength between the two subsystems. The two-photon parameter driven term $\Omega(\hat{a}^{2}+\hat{a}^{{\dagger}2})$ can be implemented by Kerr nonlinear Hamiltonian PRL120-093601 ; PRL120-093602 ; PRA444704 ; NPJQI318 ; PRL124073602 ; Nature58413205 .
In the regime of weak coupling, through the rotating wave approximation, the quantum Rabi model can be effectively described by the Jaynes-Cummings model JPB38S551 , which has been investigated widely in circuit QED system PRX2021007 , cavity QED system RMP73565 , and trapped ions PRL118073001 ; PRL128160504 . When the coupling strength approaches or surpasses the frequency scales of the cavity mode and the qubit, the rotating wave approximation no longer holds and remarkable phenomena emerge NRP1-19 . We will examine the phase diagram with respect to the two-photon-driven strength and the coupling rate between the subsystems. Then the dynamical behavior to enhance the entanglement by reinforcement learning method is investigated.
@@ III The Phase Diagram
where $\rho_{M}$ is the density matrix for the ground state of dimension-$M$ eigenspace and the integer $q$ represents the dimension difference between the two density matrices. 0-padding is used at the higher truncation margin of the density matrix $\rho_{M}$ to compensate for the truncation difference between $\rho_{M}$ and $\rho_{M+q}$. When calculating the logarithm of the matrix elements, 0 matrix elements will be replaced by 1. This metric is used to quantify the difference of $\rho_{M}$ and $\rho_{M+q}$. A smaller value of $KL_{M}^{q}$ indicates a larger similarity between the two states. As shown in Fig. 1 $(a)$ with $K/\delta_{c}=0$, $KL_{M}^{1}$ diverges when $\Omega/\delta_{c}>0.5$. This divergence behavior of $KL_{M}^{1}$ occurs with the continuous change of $\Omega/\delta_{c}$, so
@@
This phase transition can also be indicated by the energy gap between the first excited state and the ground state of the Hamiltonian (1) as shown in Fig. 1 $(b)$. As shown in $(b_{1})$, the energy spectrum exhibits a tendency towards degeneracy at $\Omega_{c}$. These results coincide with the that in Fig. 1 $(a)$. The Kerr nonlinearity $K$ in the Hamiltonian (1) can be seen as a kind of photon-dependent photon number shift PRL120-093601 ; PRL120-093602 ; PRA444704 ; NPJQI318 ; PRL124073602 ; Nature58413205 . The influence of this Kerr nonlinearity on $E_{q}-E_{0}$ can be seen in Fig. 1($b_{2}$). With increasing of $K/\delta_{c}$, the energy gap $E_{1}-E_{0}$ extends. In this work, we focus on the dynamical character while enhancing entanglement by machine learning with $K/\delta_{c}=0$.
Figure 2: $(a)$ The average of evenly distributed 10000 sampling points of the Wigner function versus $g/\delta_{c}$ and $\Omega/\delta_{c}$ for the cavity ground state. $(a_{1})$ $(a_{3})$ The distribution of $|\rho|^{2}$ for the cavity ground states when the coupling strength $g/\delta_{c}=0.3$ and $\Omega/\delta_{c}=0.3,0.5,0.503$ with the corresponding Wigner functions, respectively.
@@ IV Enhance Entanglement by Reinforcement Learning
Taking a cue from control strategies by Lyapunov control and machine learning PLA425127874 ; andp202400056 , we propose a scheme harnessing an RL agent to design temporal control field to enhance the entanglement and check the influence of system parameters. Drawing inspiration from time-dependent parameter process PRL120-093601 ; PRL120-093602 ; PRA444704 ; NPJQI318 ; PRL124073602 ; Nature58413205 , in this control proposal, the Hamiltonian for the controlled quantum system reads:
@@
here $\Omega(t)=\Omega_{max}f(t)$ with $f(t)$ is the control sequence designed by the RL agent, and $\Omega_{max}$ is the amplitude. In this control process, $[\hat{a}^{{\dagger}}\hat{a}+\hat{\sigma}_{z}+(\hat{a}+\hat{a}^{{\dagger}})(\hat{\sigma}_{-}+\hat{\sigma}_{+}),\hat{a}^{2}+\hat{a}^{{\dagger}2}]\neq 0$ ensures the effectiveness of the control fields. The workflow for this control scheme is illustrated in Fig. 3. The control field $\Omega(t)$ acts as the sequence of the actions and the increase of $E^{T}$ is positively related with the reward function in the reinforcement learning.
@@ IV.1 Reinforcement Learning
Starting with no prior knowledge about the system under control, RL uses the trial-and-error paradigm to iteratively learn the mapping between actions and states that maximizes the accumulated reward over time. An appropriate rewards-evaluation rule that favors particular state-action mappings which can enhance the control performance. Decision-making executed by the agent entails selecting actions $a_{t}=\pi(s_{t})$ that manipulates the system changing from a state (not referring to quantum states, but a set of expectations of operators of the controlled system) $s_{t}$ to $s_{t+1}$, with $\pi$ denoting the policy being learned Murphy2012 ; Sutton2018 .
@@
Figure 3: The workflow of using reinforcement learning agent to enhance the entanglement for the two-photon-driven Rabi model.
@@ IV.3 Entanglement dynamics in open conditions
where $\mathcal{L}_{\hat{a}}$ and $\mathcal{L}_{\hat{\sigma}_{-}}$ are Liouvillian superoperators describing the decoherence of the cavity field and qubit PRA84043832 . They read $\mathcal{L}_{\hat{\chi}}\rho(t)=\sum_{j,k>j}\Gamma^{jk}_{\hat{\chi}}\bar{n}(\Delta_{kj},T)\mathcal{D}[|k\rangle\langle j|]\rho(t)+\sum_{j,k>j}\Gamma^{jk}_{\hat{\chi}}(1+\bar{n}(\Delta_{kj},T))\mathcal{D}[|j\rangle\langle k|]\rho(t)$ for $\hat{\chi}=\hat{a}(\hat{\sigma}_{-})$ represents the cavity field (qubit), $\mathcal{D}[\mathcal{O}]\rho(t)$ = $\frac{1}{2}(2\mathcal{O}\rho(t)\mathcal{O}^{\dagger}-\rho(t)\mathcal{O}^{\dagger}\mathcal{O}-\mathcal{O}^{\dagger}\mathcal{O}\rho(t))$. The relaxation coefficients $\Gamma^{jk}_{\hat{\chi}}=2\pi d(\Delta_{kj})\alpha^{2}_{\hat{\chi}}(\Delta_{kj})|C^{\hat{\chi}}_{jk}|^{2}$ with $d(\Delta_{kj})$ being the spectral density of the baths, $\alpha_{\hat{\chi}}(\Delta_{kj})$ denoting the system-bath coupling strength, $\Delta_{kj}=\omega_{k}-\omega_{j}$, and $C_{jk}^{\hat{\chi}}=-i\langle j|(\hat{\chi}-\hat{\chi}^{\dagger})|k\rangle$ $\bar{n}(\Delta_{kj},T)$ ($\bar{n}$ for short) is the thermal population at frequency $\Delta_{kj}$ with environment temperature $T$. When considering a cavity coupling to the momentum quadrature of a field in one-dimension waveguides, the spectral density $d(\Delta_{kj})$ is constant and $\alpha_{\hat{\chi}}^{2}(\Delta_{kj})\propto\Delta_{kj}$. Then the relaxation coefficients reduce to $\Gamma^{jk}_{\hat{\chi}}=\gamma_{\hat{\chi}}\,(\Delta_{kj}/\omega_{0})\,|C^{\hat{\chi}}_{jk}|^{2}$ where $\gamma_{\hat{\chi}}$ is the damping rate. These assumptions can be realized in circuit-QED. The influence of dephasing and Lamb shifts in current experiments can be deemed negligible as they do not exert significant influence PRL118073001 ; PRL128160504 .
@@
In this study, we put forward a conjecture that the dynamics of the entanglement can reflect the phase transition for different initial states. We first check the dynamics of the entanglement versus the constant two-photon-driven strength when $g/\delta_{c}=0.3$ in Fig. 4. As shown in Fig. 4 ($a$), the strength and duration time of the entanglement for the ground state decreases as $\Omega/\delta_{c}$ increases. This character coincides with the results in Fig. 1 ($c$). This means constant two-photon process has a negative effect on the entanglement. While the initial state $\rho(0)$ = $|\alpha=1,\downarrow\rangle\langle\alpha=1,\downarrow|$, where $|\alpha=1\rangle$ is the coherent state with amplitude 1, the entanglement exhibits more fluctuations versus time when $\Omega_{max}/\delta_{c}$ is below the phase critical point $\Omega_{c}$. Comparing Fig. 4 ($a$) and ($b$), the entanglement increases more obviously when the initial state is the non-entangled product state than the entangled ground state.
Figure 5: (a) The dynamics of the entanglement indicated by $E^{T}$ versus the training epoches. (b) The corresponding square control pulses $\Omega(t)$ ($\Omega_{max}/\delta_{c}$ = 0.3) from the top view, with values shown in the color bar. $\rho(0)=\rho_{0}$ is the ground state of the Hamiltonian (1), $\delta_{q}/\delta_{c}$ = 1, $g/\delta_{c}$ = 0.3, $K/\delta_{c}$ = 0, and dissipation rates $\gamma_{\hat{a}}/\delta_{c}$ = $\gamma_{\hat{\sigma}_{-}}/\delta_{c}$ = 0.01.

### v4

@@ Abstract
A control scheme is proposed that leverages reinforcement learning to enhance entanglement by modulating the two-photon-driven amplitude in a Rabi model. The quantum phase diagram versus the two-photon-driven amplitude and the coupling between the cavity field and the atom in the Rabi model is indicated by the energy spectrum of the hybrid system, the witness of entanglement, second order correlation, and negativity of Wigner function. From a dynamical perspective, the behavior of entanglement can reflect the phase transition and the reinforcement learning agent is employed to produce temporal sequences of control pulses to enhance the entanglement in the presence of dissipation. The entanglement can be enhanced in different parameter regimes and the control scheme exhibits robustness against dissipation. The replaceability of the controlled system and the reinforcement learning module demonstrates the generalization of this scheme. This research paves the way of positively enhancing quantum resources in non-equilibrium systems.
@@ I Introduction
The interaction between a two-level system and an electromagnetic field, fully described by the quantum Rabi model PR51652 ; RPP691325 is a platform which has widespread utility ranging from the study of fundamental physics Haroche2006 to the cutting edge of quantum information RMP91025005 . In the regime of cavity quantum electrodynamics (QED), when the cavity-qubit coupling rate is strong enough to dominate dissipation rates or reaches the domain of transition frequencies of the cavity or qubit, the counter-rotating term cannot be ignored, and the system is best described by the quantum Rabi model RMP91025005 ; NRP1-19 ; PRL120-183601 . Once the coupling is strong enough, a wide range of interesting phenomena occurs, including innovative quantum phase transitions in the single-qubit-single-oscillator model when the frequency of the oscillator is much lower than that of the qubit PRA87-013826 ; PRL115180404 , generation of pure-cat-state PRA81042311 ; PRA96043834 , strongly entangled states, and novel steady-state squeezing PRL131223604 . To date, in specially designed architectures, the qubit-cavity coupling rate can reach the considerable fraction of the cavity transition frequency NP6772 . By inductively coupling a flux qubit and an LC oscillator via Josephson junctions, ultrastrong-coupling regime with unconventional transition spectra has been realized  NP1344 . Quantum simulation in circuit QED provides experimental access to the physics of the ultrastrong-and deep-strong coupling regimes of an effective quantum Rabi model PRX2021007 ; NC8779 . The strong coupling limit of cavity QED has been demonstrated in a solid-state system by the vacuum Rabi mode splitting and the corresponding avoided crossings Nature431162 .
@@
Machine-learning techniques are emerging as effective tools in exploring physics Murphy2012 ; Sutton2018 ; Murphy2018 ; PRX8031084 , and among them, reinforcement learning (RL) has the potential to find optimal control fields to engineer system dynamics, given that the evolution of these systems is governed by specific differential equations. Time-varying control proposals have been employed to produce spin squeezing in a collective spin systems PRA103032601 ; AdP5362400056 . These works provide the fundamental model to generate specific quantum states by reinforcement learning in this work. Reinforcement learning can successfully steer the quantum Kapitza oscillator to a Floquet-engineered stable inverted position amidst noise, with applications in quantum information and quantum opticsPRB98224305 . The application of RL to quantum control for a non-Markovian open quantum system preparing the superposition of two excited states highlights the potential of RL for quantum control in complex systems PRA109013104 .
@@ II The model
where $\hat{a}^{\dagger}$($\hat{a})$ is the creation (annihilation) operator for the bosonic cavity mode and $\hat{\sigma}_{z}$ is the Pauli $z$-basis operator with commute relation $[\hat{\sigma}_{+},\hat{\sigma}_{-}]=\hat{\sigma}_{z}$. $\delta_{c(q)}=\omega_{c(q)}-\omega_{p}$, where $\omega_{c}$, $\omega_{q}$ and $\omega_{p}$ are the frequencies of the cavity, qubit and parametrical driving, respectively. The parameter $g$ is the coupling strength between the two subsystems. The two-photon-driven term $\Omega(\hat{a}^{2}+\hat{a}^{{\dagger}2})$ can be implemented by Kerr nonlinear Hamiltonian PRL120-093601 ; PRL120-093602 ; PRA444704 ; NPJQI318 ; PRL124073602 ; Nature58413205 .
[…]
@@ III The Phase Diagram
where $\rho_{M}$ is the density matrix for the ground state of dimension-$M$ eigenspace and the integer $q$ represents the dimension difference between the two density matrices. 0-padding is used at the higher truncation margin of the density matrix $\rho_{M}$ to compensate for the truncation difference between $\rho_{M}$ and $\rho_{M+q}$. When calculating the logarithm of the matrix elements, 0 matrix elements will be replaced by 1. This metric is used to quantify the difference between $\rho_{M}$ and $\rho_{M+q}$. A smaller value of $KL_{M}^{q}$ indicates a larger similarity between the two states. As shown in Fig. 1 $(a)$ with $K/\delta_{c}=0$, $KL_{M}^{1}$ diverges when $\Omega/\delta_{c}>0.5$. This divergence behavior of $KL_{M}^{1}$ occurs with the continuous change of $\Omega/\delta_{c}$, so
@@
This phase transition can also be indicated by the energy gap between the first excited state and the ground state of the Hamiltonian (1) as shown in Fig. 1 $(b)$. As shown in $(b_{1})$, the energy spectrum exhibits a tendency towards degeneracy at $\Omega_{c}$. These results coincide with the that in Fig. 1 $(a)$. The Kerr nonlinearity $K$ in the Hamiltonian (1) can be seen as a kind of photon-dependent photon number shift PRL120-093601 ; PRL120-093602 ; PRA444704 ; NPJQI318 ; PRL124073602 ; Nature58413205 . The influence of this Kerr nonlinearity on $E_{1}-E_{0}$ can be seen in Fig. 1($b_{2}$). With increasing of $K/\delta_{c}$, the region of energy gap $E_{1}-E_{0}=0$ extends. In this work, we focus on the dynamical character while enhancing entanglement by machine learning with $K/\delta_{c}=0$.
Figure 2: $(a)$ The average of evenly distributed 10000 sampling points of the Wigner function versus $g/\delta_{c}$ and $\Omega/\delta_{c}$ for the cavity ground state. $(a_{1})$ $(a_{3})$ The distribution of $|\rho|^{2}$ for the cavity ground states when the coupling strength $g/\delta_{c}=0.3$ and $\Omega/\delta_{c}=0.1,0.5,0.503$ with the corresponding Wigner functions, respectively.
@@ IV Enhance Entanglement by Reinforcement Learning
Taking a cue from control strategies by Lyapunov control and machine learning PLA425127874 ; AdP5362400056 , we propose a scheme harnessing an RL agent to design temporal control field to enhance the entanglement and check the influence of system parameters. Drawing inspiration from time-dependent parameter process PRL120-093601 ; PRL120-093602 ; PRA444704 ; NPJQI318 ; PRL124073602 ; Nature58413205 , in this control proposal, the Hamiltonian for the controlled quantum system reads:
@@
here $\Omega(t)=\Omega_{max}f(t)$ with $f(t)$ is the control sequence designed by the RL agent, and $\Omega_{max}$ is the amplitude. In this control process, $[\hat{a}^{{\dagger}}\hat{a}+\hat{\sigma}_{z}+(\hat{a}+\hat{a}^{{\dagger}})(\hat{\sigma}_{-}+\hat{\sigma}_{+}),\hat{a}^{2}+\hat{a}^{{\dagger}2}]\neq 0$ ensures the effectiveness of the control fields. The workflow for this control scheme is illustrated in Fig. 3. The control field $\Omega(t)$ acts as the sequence of the actions and the increase of $E^{T}$ is positively related with the reward function in the reinforcement learning module.
@@ IV.1 Reinforcement Learning
Starting with no prior knowledge about the system under control, RL uses the trial-and-error paradigm to iteratively learn the mapping between actions and states that maximizes the accumulated reward over time. An appropriate rewards-evaluation rule which favors particular state-action mappings can enhance the control performance. Decision-making executed by the agent entails selecting actions $a_{t}=\pi(s_{t})$ that manipulate the system changing from a state (not referring to quantum states, but a set of expectations of operators of the controlled system) $s_{t}$ to $s_{t+1}$, with $\pi$ denoting the policy being learned Murphy2012 ; Sutton2018 .
@@
Figure 3: The workflow of using reinforcement learning agent to enhance the entanglement for the two-photon-driven Rabi model. The hyper parameters for the DQN agent are set as follows: Batchsize = 128. The discount factor is set to 0.99. The starting value of $\varepsilon$ for the greedy strategy is set to 0.9 and the final value of $\varepsilon$ is 0.05, with exponential decay 1000. The update rate of the target network is set to 0.005. The learning rate of the AdamW optimizer is set to $1\times 10^{-4}$. The neural network architecture used by the DQN agent is defined as follows: The network takes as input the observable properties of the quantum system is set to 6. Hidden Layers include layer1: fully connected layer has input size = 128 and output size = 128 with ReLU activation function. layer2: fully connected layer has input size = 128 and output size = 512, also with ReLU activation function. layer3: fully connected layer with input size = 512 and output size = 256, also with ReLU activation function. Output Layer: fully connected layer with input size = 256 and output size = 3 correspond to the possible actions. The output layer directly returns the Q-values without using an activation function.
@@ IV.3 Entanglement dynamics in open conditions
where $\mathcal{L}_{\hat{a}}$ and $\mathcal{L}_{\hat{\sigma}_{-}}$ are Liouvillian superoperators describing the decoherence of the cavity field and qubit PRA84043832 . They read $\mathcal{L}_{\hat{\chi}}\rho(t)=\sum_{j,k>j}\Gamma^{jk}_{\hat{\chi}}\bar{n}(\Delta_{kj},T)\mathcal{D}[|k\rangle\langle j|]\rho(t)+\sum_{j,k>j}\Gamma^{jk}_{\hat{\chi}}(1+\bar{n}(\Delta_{kj},T))\mathcal{D}[|j\rangle\langle k|]\rho(t)$ for $\hat{\chi}=\hat{a}(\hat{\sigma}_{-})$ represents the cavity field (qubit), $\mathcal{D}[\mathcal{O}]\rho(t)$ = $\frac{1}{2}(2\mathcal{O}\rho(t)\mathcal{O}^{\dagger}-\rho(t)\mathcal{O}^{\dagger}\mathcal{O}-\mathcal{O}^{\dagger}\mathcal{O}\rho(t))$. The relaxation coefficients $\Gamma^{jk}_{\hat{\chi}}=2\pi d(\Delta_{kj})\alpha^{2}_{\hat{\chi}}(\Delta_{kj})|C^{\hat{\chi}}_{jk}|^{2}$ with $d(\Delta_{kj})$ being the spectral density of the baths, $\alpha_{\hat{\chi}}(\Delta_{kj})$ denoting the system-bath coupling strength, while $\Delta_{kj}=\omega_{k}-\omega_{j}$, and $C_{jk}^{\hat{\chi}}=-i\langle j|(\hat{\chi}-\hat{\chi}^{\dagger})|k\rangle$ $\bar{n}(\Delta_{kj},T)$ ($\bar{n}$ for short) is the thermal population at frequency $\Delta_{kj}$ with environment temperature $T$. When considering a cavity coupling to the momentum quadrature of a field in one-dimension waveguides, the spectral density $d(\Delta_{kj})$ is constant and $\alpha_{\hat{\chi}}^{2}(\Delta_{kj})\propto\Delta_{kj}$. Then the relaxation coefficients read $\Gamma^{jk}_{\hat{\chi}}=\gamma_{\hat{\chi}}\,(\Delta_{kj}/\omega_{0})\,|C^{\hat{\chi}}_{jk}|^{2}$ where $\gamma_{\hat{\chi}}$ is the damping rate. These assumptions can be realized in circuit-QED. The influence of dephasing and Lamb shifts in current experiments can be deemed negligible as they do not exert significant influence PRL118073001 ; PRL128160504 .
@@
In this study, we put forward a conjecture that the dynamics of the entanglement can reflect the phase transition for different initial states. We first check the dynamics of the entanglement versus the constant two-photon-driven strength when $g/\delta_{c}=0.3$ in Fig. 4. As shown in Fig. 4 ($a$), the strength and duration time of the entanglement for the ground state decreases as $\Omega/\delta_{c}$ increases. This character coincides with the results in Fig. 1 ($c$). This means the two-photon process has a negative effect on the entanglement. While the initial state $\rho(0)$ = $|\alpha=1,\downarrow\rangle\langle\alpha=1,\downarrow|$, where $|\alpha=1\rangle$ is the coherent state with amplitude 1, the entanglement exhibits more fluctuations versus time when $\Omega_{max}/\delta_{c}$ is below the phase critical point $\Omega_{c}$. Comparing Fig. 4 ($a$) and ($b$), the entanglement increases more obviously when the initial state is the non-entangled product state than the entangled ground state.
Figure 5: (a) Evolution of entanglement witness $E^{T}$ versus $\Omega_{max}/\delta_{c}$ for $\rho(0)=\rho_{0}$ (the ground state of the Hamiltonian (1)), and (b) $\rho(0)$ = $|\alpha=1,\downarrow\rangle\langle\alpha=1,\downarrow|$. $\delta_{q}/\delta_{c}$ = 1, $g/\delta_{c}$ = 0.3, $K/\delta_{c}$ = 0, and dissipation rates $\gamma_{\hat{a}}/\delta_{c}$ = $\gamma_{\hat{\sigma}_{-}}/\delta_{c}$ = 0.01.